Maize and Blue Co. is a multinational enterprise (MNE) that manufactures many different products as parts for the aerospace industry. The new controller, Mary, has been asked by the president of the company to provide accounting data for decisions she needs to make. Each question is worth 1 point.  

Special Order: Suppose Wolverine Corp. approached your company, Maize and Blue Co., with a special order. Wolverine Corp. wishes to purchase 63,000 "warning" decals for its electrical panels and offers to pay you $0.86 per decal. Your company has enough capacity to handle the special order and its total production cost is $0.56 per decal, as follows:

Variable costs:
Direct materials $0.13
Direct labor $0.10
Variable overhead $0.13
Fixed overhead $0.20
Total cost $0.56

1. What are the total relevant costs if Maize and Blue accepts the special order?

2. What is the increase or decrease in net operating income from the order if Maize and Blue accepts the special order? Enter an increase as a positive number and a decrease as a negative number.

3. Discontinuing a Product Line: Maize and Blue is considering discontinuing one of its product lines. During the past year, the product line's income statement showed the following:

Sales revenue $7,389,000
Less: cost of goods sold $6,500,000
Gross profit $889,000
Less: operating expenses $1,600,000
Operating income (loss) $-711,000

Fixed manufacturing overhead costs account for 10% of the cost of goods, while only 30% of the operating expenses are fixed. Since the product line is just one of Maize and Blue's products, only $740,000 of total fixed expenses (the majority of which is advertising) will be eliminated if the product line is discontinued. If the company decides to discontinue the product line, how much will the company's net operating income increase or decrease? Enter an increase as a positive number and a decrease as a negative number.

Answer questions 4 and 5 with the following information:

Maize and Blue packages small LED lights for plane instrument panels. Cost data for this packaging process are as follows:  

Unit Cost
Packaging materials (e.g., boxes, bubble-wrap) $2.18
Packaging direct labor $0.67
Indirect materials (e.g., tape, labels) $0.42
Packaging supervision (variable) $0.40
Other fixed manufacturing overhead $1.46
Total packaging cost $5.13



An outside supplier has offered to do all the packaging for a price of $4 per unit for all packaging-related activities if Maize and Blue signs a one-year contract for a minimum of 143,400 units produced each year. Maize and Blue could use the factory space now occupied by the packaging process to expand production to another product line. This expansion is expected to generate an additional $155,000 in profit per year.  

4. Make or Buy: What are the total relevant costs of continuing to package the products internally; that is, which of the annual costs is avoidable if Maize and Blue outsources the packaging process?

5. Make or Buy: What is the relevant cost of outsourcing the packaging process considering the profit from expanding production of another product?
Answer #1

1. Total relevant costs are the costs which are incurred for the special order.

Since the company has spare capacity, fixed costs will not be incurred for this order and are not relevant

Total relevant cost = (0.13+0.10+0.13)*63,000 = $22,680

2.Increase in operating income = (0.86-0.36)*63,000

=$31,500

3. Increase in operating Income =savings in fixed costs - contribution margin lost

= 740,000 -(7,389,000-6,500,000*90% - 1,600,000*70%)

= $321,000

4. Fixed costs will continue to occur irrespective of the decision

Hence, total avoidable cost will be the variable costs

= (2.18+0.67+0.42+0.40)*143,400

=$526,278

5.Relevant cost = 4*143,400 - 155,000

=$418,600

Hence, buying is better